OpenStudy (anonymous):

A farmer has 180 feet of fence to create a rectangular pen next to his barn. No fence is needed on the barn side. Find the dimensions for the maximum area for the pen.

OpenStudy (anonymous):

so far i have 180=2x+y 180 -2x = y Find max A: A=xy A=x(180-2x) A=180x-2x^2

OpenStudy (anonymous):